OpenStudy (anonymous):

Can someone solve this using elimination please? .40x+.35y=56 .30x+.25y=41

OpenStudy (anonymous):

Separately?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

I thought I knew but I dont. All I have is: .40x+.35y=56 -.35y -.35y .40x=56-.35y -56 But that would leave you with .40x=.35y What do you do then, sorry I thought I could help.

OpenStudy (anonymous):

(70,80) i think

OpenStudy (anonymous):

Thank you Nickie_S and satellite73 for helping me :)

OpenStudy (anonymous):

best method if you want to do it for yourself is to multiply both equations by 100 to clear the annoying decimals and start with \[40x+35y=5600\] \[30x+25y=4100\] then maybe divide everything by 5 to get \[8x+7y=1120\] \[6x+5y=224\] and now you have numbers that are more reasonable

OpenStudy (anonymous):

Thank you so much satellite73!!!! <3333333

OpenStudy (anonymous):

yw. still did not solve it though. you have to multiply first equation by -5 and second by 7 then add to get the answer. pretty sure my answer is right though

OpenStudy (anonymous):

Oh i solved it :)

OpenStudy (anonymous):

And yes, the answer is correct